### v3

# MIRACLE II: Unveiling the multi-phase gas interplay in the circumnuclear region of NGC 1365 via multi-cloud modeling

11email: matteo.ceci@unifi.it 22institutetext: INAF - Osservatorio Astrofisico di Arcetri, Largo E. Fermi 5, I-50125 Firenze, Italy 33institutetext: Scuola Normale Superiore, Piazza dei Cavalieri 7, 56126 Pisa, Italy 44institutetext: Instituto de Radioastronomía y Astrofísica, Universidad Nacional Autónoma de México, Morelia, Michoacán 58089, Mexico 55institutetext: ESO, Karl-Schwarzschild-Str. 2, 85748 Garching bei München, Germany 66institutetext: Instituto de Astrofísica de Canarias, 38205 La Laguna, Tenerife, Spain 77institutetext: Departamento de Astrofísica, Universidad de La Laguna, 38206 La Laguna, Tenerife, Spain 88institutetext: Institute of Physics, GALSPEC laboratory, EPFL, Observatory of Sauverny, Chemin Pegasi 51, 1290 Versoix, Switzerland 99institutetext: AURA for ESA, Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 1010institutetext: INAF, Osservatorio Astronomico di Trieste, Via Tiepolo 11, I-34131 Trieste, Italy 1111institutetext: IFPU - Institute for Fundamental Physics of the Universe, via Beirut 2, I-34151 Trieste, Italy 1212institutetext: University of Trento, Via Sommarive 14, Trento, I-38123, Italy 1313institutetext: Max-Planck-Institut für Extraterrestrische Physik (MPE), Gießenbachstr. 1, D-85748 Garching, Germany 1414institutetext: Observatorio Astronómico Nacional, C/ Alfonso XII 3, 28014 Madrid, Spain 1515institutetext: Dipartimento di Fisica e Astronomia, Alma Mater Studiorum, Università degli Studi di Bologna, Via Gobetti 93/2, 40129 Bologna, Italy 1616institutetext: INAF–Osservatorio di Astrofisica e Scienza dello Spazio di Bologna, Via Gobetti 93/3, 40129 Bologna, Italy
MIRACLE II: Unveiling the multi-phase gas interplay in the circumnuclear region of NGC 1365 via multi-cloud modeling
We present a multi-phase analysis of the gas in the circumnuclear region ( 0.90.9 kpc2) of the nearby barred Seyfert 1.8 galaxy NGC 1365, observed as part of the Mid-IR Activity of Circumnuclear Line Emission (MIRACLE) program. Specifically, we combined spatially resolved spectroscopic data from JWST/MIRI, VLT/MUSE and ALMA to provide a multi-phase characterization of the ionized atomic and the warm and cold molecular gas phases. MIRI data enabled the detection of more than 40 mid-IR emission lines from ionized or warm molecular gas. Moment maps show that both cold and warm molecular gas trace the circumnuclear ring, following the rotation of the stellar disk. The ionized gas exhibits flux distributions and kinematics varying depending on the ionization potential (IP). Low-IP species ( 25 eV) mainly trace the rotating disk, while higher-IP species (up to 120 eV) trace the outflowing gas. Both [O III] and [Ne V] trace the nuclear outflow cone towards the southeast. In addition, the [Ne V] line traces the counter-cone of the outflow to the northwest, which is obscured in the optical at these circumnuclear scales and thus undetected in [O III]. Unlike optical diagnostics, spatially resolved mid-IR diagnostics reveal the key role of the active galactic nuclei (AGN) as the source of gas ionization in the central region. We derive the electron density from the [Ne V]/[Ne V] line ratio, finding a median value of (750440) cm-3, consistent with previous estimates obtained from the optical [S II] doublet. Lastly, we apply, for the first time, a fully self-consistent combination of state-of-the-art photoionization and kinematic models (HOMERUN+) to estimate the intrinsic physical outflow properties, kinematics, and energetics - overcoming the limitations of classical methods based on oversimplified assumptions. Exploiting the unprecedented synergy between JWST/MIRI and VLT/MUSE, HOMERUN allows us to simultaneously reproduce the fluxes of over 60 emission lines spanning from the optical to the mid-IR. This unique approach enables us to disentangle the physical conditions of AGN- and star formation-dominated components and robustly estimate the mass of the outflowing gas and other physical properties.
Key Words.:
galaxies: Seyfert - galaxies: ISM - galaxies: active - ISM: kinematics and dynamics - ISM: jets and outflows1 Introduction
Galactic outflows are known to influence the evolution of their host galaxies by removing gas, enriching the circumngalactic medium, and suppressing star formation in the galaxy (Veilleux et al., 2005; Fabian, 2012; Cresci & Maiolino, 2018; Harrison & Ramos Almeida, 2024). While powerful outflows are commonly observed in luminous active galactic nuclei (AGN) at redshifts (e.g., Cano-Díaz et al., 2012; Fiore et al., 2017), similar signatures are also detected in local, less-powerful AGN host galaxies, where high spatial resolution enabled by their vicinity allows for detailed studies of their physical and kinematic properties (Cresci et al., 2015; Venturi et al., 2017, 2018; Cicone et al., 2018; Mingozzi et al., 2019; Fluetsch et al., 2019; Lutz et al., 2020). These local outflows span multiple gas phases, from ionized atomic to molecular, and show broad, blue-shifted components in emission and/or absorption lines, consistent with being driven by AGN activity.
Ionized outflows, often traced by strong optical emission lines such as [O III], exhibit velocities ranging from several hundreds to a few thousands km s-1. They are composed of warm gas (T = K, ne=cm-3) that expands from the nuclear region out to galactic scales, potentially expelling material from their host galaxy (e.g., Harrison et al., 2014, 2016; Woo et al., 2016; Fiore et al., 2017; Venturi et al., 2018; Förster Schreiber et al., 2019; Kakkad et al., 2020; Venturi et al., 2021; Cresci et al., 2023; Speranza et al., 2024; Marconcini et al., 2025b). However, the impact of ionized winds on star formation (SF) remains debated (e.g., Harrison et al., 2016; Harrison & Ramos Almeida, 2024), as their coupling efficiency with the ambient interstellar medium (ISM) appears to be low. Indeed, the energy and momentum carried by the ionized gas phase are typically sub-dominant when compared to those of other gas phases (e.g. Zubovas & Bourne, 2017; Combes, 2017; Fluetsch et al., 2019, 2021; Mulcahey et al., 2022; Belli et al., 2024). In the context of these findings, the ionized gas component may only trace the larger scale, less massive regions of multi-phase outflows, which are dominated in mass by colder molecular or atomic gas on less extended scales (e.g. Ramos Almeida et al., 2022; Audibert et al., 2023; Venturi et al., 2023).
Cold molecular outflows represent a colder phase of gas (T = K, n cm-3), frequently observed via CO emission lines. Compared to the ionized phase, CO-traced outflows are typically slower, with velocities of a few hundred km s-1, but often carry substantial mass, suggesting they may play a crucial role in regulating SF (Feruglio et al., 2010; Fiore et al., 2017; García-Burillo et al., 2019; Lutz et al., 2020; Fluetsch et al., 2019, 2021). In addition to the cold component, warm molecular outflows (T = K) are typically traced by H2 roto-vibrational transitions in the near- and mid-infrared (IR). H2 emission is particularly prominent in shocked regions, making it a sensitive tracer of warm molecular gas entrained in outflows (Hill & Zakamska, 2014; Richings & Faucher-Giguère, 2018a, b; Riffel et al., 2020; Wright et al., 2023).
NGC 1365, the subject of this work, is a great laboratory to study in detail multi-phase AGN feedback. This dusty barred spiral galaxy (SB(s)b; de Vaucouleurs et al. 1991) is located in the Fornax cluster (Jones & Jones, 1980) at a distance of 19.57 Mpc (Jacobs et al., 2009; Anand et al., 2021) (1″ 95 pc). With a redshift of z = 0.005457 (Bureau et al., 1996), NGC 1365 is classified as a Seyfert 1.8 galaxy (Véron-Cetty & Véron, 2006), with signature of both AGN activity and SF in its central regions.
The AGN-driven ionized outflow in NGC 1365 is observed via extended biconical [O III] emission spanning 2.5 kpc, with line-of-sight velocities up to 170 km s-1 (Venturi et al., 2018). The southeastern (SE) cone, approaching the observer, and the receding northwestern (NW) cone - partially obscured by the galactic disk - are dominated by AGN ionization as traced by diagnostic diagrams (Baldwin et al., 1981; Veilleux & Osterbrock, 1987; Kewley et al., 2006). High-resolution X-ray observations revealed a fast, highly ionized blue-shifted nuclear (spatially unresolved) wind in absorption with velocities of 3000 km s-1, further supporting the scenario of an AGN-driven wind (Risaliti et al., 2005; Braito et al., 2014).
NGC 1365 has a SF rate of 16.9 M⊙ yr-1 (Lee et al., 2022), mainly within a 2 k radius circumnuclear ring identified across optical, radio, and IR wavelengths (Kristen et al., 1997; Forbes & Norris, 1998; Alonso-Herrero et al., 2012). The ring is associated with the inner Lindblad resonance (Lindblad et al., 1996) and shows non-circular bar-driven gas motions (Teuben et al., 1986; Sanchez, 2009), associated with SF-driven ionization (Sharp & Bland-Hawthorn, 2010; Venturi et al., 2018) and large reservoirs of molecular gas ( ; Sakamoto et al., 2007; Gao et al., 2021). Observational campaigns from the PHANGS survey (“Physics at High Angular resolution in Nearby GalaxieS”, Lee et al. 2023) conducted with the James Webb Space Telescope (JWST) and complemented by the Atacama Large Millimeter/submillimeter Array (ALMA; Wootten & Thompson 2009) have revealed that young clusters embedded in the ring drive both ionization of the surrounding medium and localized modifications to the molecular gas phase, where feedback-driven heating and dissociation decrease CO excitation and increase [C I]/CO abundance ratios in their immediate vicinity (Liu et al., 2023; Schinnerer et al., 2023).
Despite hints for the presence of a nuclear radio jet in NGC 1365 (Sandqvist et al., 1995), later studies by Stevens et al. (1999) found no conclusive evidence for its presence. Instead, radio emission is primarily attributed to the elongated star-forming ring rather than a jet.
JWST is transforming our understanding of the outflow properties in the local Universe by providing an unprecedented view of the nuclear dusty regions of AGN (e.g. Hermosa Muñoz et al., 2024a; Perna et al., 2024; Zhang et al., 2024; García-Bernete et al., 2024; Davies et al., 2024; Ulivi et al., 2025; Ceci et al., 2025). This revolution is largely driven by the unique capabilities of the JWST Mid-Infrared Instrument (MIRI; Wright et al., 2015, 2023), including its high spatial resolution across a wide range of wavelengths (from 5 m to 28 m), enabling detailed studies of the energetics and kinematics of gas flows in obscured nuclear regions of AGN. The wide IR spectral range of JWST spanning many atomic and molecular gas transitions, combined with its spatial and spectral resolution, provides essential insights into the multi-phase circumnuclear medium. These unprecedented capabilities enable multi-phase, co-spatial analyses when combined with complementary data from ground-based facilities, such as the Multi Unit Spectroscopic Explorer (MUSE) at the ESO Very Large Telescope (VLT; Bacon et al. 2010) and ALMA, transforming our understanding of feedback mechanisms in local AGN.
For NGC 1365, we combine data from the Medium-Resolution Spectrometer (MRS; Rieke et al. 2015; Labiano et al. 2021) of MIRI with integral-field optical spectroscopy from VLT/MUSE and millimeter observations from ALMA. This multi-wavelength approach enables us to trace the multi-phase gas across a wide range of ionization states and densities. In particular, it allows for a detailed characterization of the kinematics, morphology, and energetics of the circumnuclear environment in NGC 1365, ultimately contributing to a deeper understanding of AGN feedback and its impact on galaxy evolution.
This paper is organized as follows. In Section 2, we introduce the observations and the spectroscopic analysis of the MIRI data. In Section 3, we discuss our results through the characterization of the ISM properties and the spatially resolved kinematics of the ionized and molecular gas in the central region of NGC 1365. Finally, in Section 4, we summarize our results. For all the maps shown in this work, north is up and east is to the left.
2 Overview of observations and data analysis
NGC 1365 was observed on 8th December 2024 as part of the “Mid-IR Activity of Circumnuclear Line Emission” (MIRACLE; JWST GO program 6138; Co-PIs: C. Marconcini and A. Feltre), aimed at tracing the mid-IR emission by exploiting MIRI/MRS data of the circumnuclear region of local AGN in the 5-28 m wavelength range. The MRS mode of MIRI covers a total wavelength range of 4.9–27.9 m, divided into four integral field units (IFUs), also referred to as channels (Ch1, Ch2, Ch3, and Ch4, hereafter), each further subdivided into three bands (SHORT, MEDIUM, and LONG). These channels cover slightly different fields of view (FoVs), from 3.2″ 3.7″ in Ch1 to 6.6″ 7.7″ in Ch4, at varying pixel size (from 0.13″ in Ch1 to 0.35″ in Ch4) and resolving powers (from 3700 to 1500; see e.g. Labiano et al., 2021; Argyriou et al., 2023). As shown in Law et al. (2023), the average FWHM of the MIRI/MRS PSF ranges from 0.4″ in Ch1 to 0.9″ in Ch4. We summarize the data reduction in Appendix A.1, referring to Marconcini et al. (2025a) for a detailed description. As a result of this process, the pipeline produced 12 datacubes, one for each sub-band.
The science observations targeted the nuclear region of NGC 1365 but were deliberately offset to include the inner part of the SE approaching outflow. As shown in the lower panels of Fig. 1, the different MIRI FoVs of each channel are not exactly centered on the nucleus, but shifted toward the SE to better capture the outflow region. This choice was motivated by the primary goal of our MIRACLE project of characterizing multi-phase gas outflows and nuclear properties. In the lower right panel of Fig. 1, we show the nuclear region of NGC 1365 in the MIRI FoV. We extracted integrated spectra from the 0.7″ radius circular apertures, and we present them in Fig. 2. The green (blue) spectrum is extracted from the circular region shown in Fig. 1 where the outflow (stellar disk) emission is dominant, as is shown by the flux maps in Fig. 3. Note that the stellar disk spectrum is contaminated by outflow emission, as indicated by the presence of high-excitation lines. We detected more than 40 emission lines, with a signal-to-noise ratio (S/N) 5, in the spectral range covered by MIRI, mostly tracing ionized gas, with an ionization potential IP - defined as the energy required to create the relevant species - ranging from a few eV to more than 100 eV (see Table LABEL:tab:_table_list_emission_lines). Additionally, we also detected two hydrogen recombination lines, namely Pf and Hu (i.e. H I (6-5) at 7.46 m and H I (7-6) at 12.37 m, respectively) and seven H2 pure-rotational transitions, from 0-0 S(7) to 0-0 S(1).
As shown in Fig. 2, the MIRI spectra also exhibit at least 15 bright PAH features in the wavelength range between 5 - 18 m, which we will investigate in a dedicated forthcoming paper. To obtain these spectra, for the first time we combined the emission from all 12 MIRI datacubes, taking into account the different pixel sizes and ensuring a proper flux conservation among different bands by applying a scaling factor to align the flux levels between adjacent bands. A detailed description of this novel procedure is presented in the Appendix B. This algorithm, written in Python, is available for download111https://github.com/matteo-ceci/JWST-MIRI_stitching_code.
In Appendix C.1, we analyzed the MIRI MRS datacubes using a customized python script to enhance the S/N with spatial smoothing, subtract the local continuum around all the brightest emission lines and then fit them spaxel by spaxel. The fitted emission lines are listed in Table LABEL:tab:_table_list_emission_lines.
The MUSE IFU observations of NGC 1365 were obtained on the 12th October 2014, under program 094.B-0321(A) (PI A. Marconi) and are part of the “Measuring AGN Under MUSE” (MAGNUM) survey (Cresci et al., 2015; Venturi et al., 2017, 2018; Mingozzi et al., 2019; Marconcini et al., 2023, 2025b). We applied the same data reduction pipeline that has been extensively applied in previous works (e.g. Venturi et al., 2018; Mingozzi et al., 2019), and which has been validated and described in detail therein. To trace the cold molecular gas, we used archival ALMA 12-m Band 7 observations (program 2016.1.00296.S, PI F.Combes; Combes et al. 2019) of the CO J= transition at 345.796GHz (rest-frame). The observation description, data reduction and emission line fitting procedure of the MUSE and ALMA data can be found in Appendices A.2-C.2 and A.3-C.3, respectively.
3 Results
3.1 Multi-phase gas kinematics
For the kinematic analysis, we focused on emission lines within MIRI Ch3 (0.2″/pixel in a 6.6″ 7.7″ FoV, corresponding to 19 pc/pixel over a 620 730 pc2 FoV) because of the good balance between spatial resolution and field coverage. Moreover, the Ch3 emission lines are among the brightest detected in our data, allowing us to present spatially resolved results representative of the ionized and molecular emission, with sufficient S/N across the entire FoV. Although in this analysis we focus on the FoV of Ch3, which also includes a fraction of disk gas illuminated by the AGN and not part of the outflow, it would be more accurate to refer to this region as the ionization cone rather than calling it outflow. However, we will assume in the following that all the gas in the ionization cone is in outflow.
Figure 3 shows the moment maps of [Ne V], [Ne II], H2 S(1), and CO(3-2), obtained from the emission line fitting routines presented in Appendices C.1 and C.3. The [Ne V] seems to follow the direction of the ionized outflow, that is, the SE-NW direction (Venturi et al. 2018), as expected from species with high-IP (see Table LABEL:tab:_table_list_emission_lines). Its velocity field is also consistent with the [O III] kinematics, which is blueshifted in the SE and redshifted in the NW. In the SE (NW) cone, we estimate gas projected velocities up to -120 km s-1 (+100 km s-1). The velocity dispersion map of [Ne V] shows a donut-shaped structure in SE direction of the nucleus (highlighted with contours in Fig. 4), at a projected distance of 2.5″ ( 240 pc), with a peak of 170 km s-1.
The lower-IP [Ne II] emission line has a different morphology with respect to the highly ionized gas traced by the [Ne V], suggesting that it traces a different gas component. Indeed, the [Ne II] morphology resembles the stellar continuum emission in the circumnuclear ring (whose ALMA contours are highlighted in Fig. 1; see also Liu et al., 2023), as we can see in Fig. 3. Moreover, the moment-0 map reveals bright clumps and lanes in the left part of the FoV that match remarkably well the features visible in H map of Venturi et al. (2018). This supports the interpretation that the [Ne II] line predominantly traces the circumnuclear ring. This includes the elongated structure 4″ eastward of the nucleus, which is part of the so-called ‘Mid-East’ region - associated with the Southern Arm, a stream of gas flowing in from the southeast along the bar - and is interpreted as material moving downstream within the circumnuclear ring (Liu et al., 2023). The fact that [Ne II] traces the disk component is also confirmed by the velocity gradient of [Ne II], whose direction, along the galaxy disk major axis, and magnitude ( km s-1) are consistent with those of the rotating stellar component (and H) observed on the same scales by Venturi et al. (2018).
The warm molecular gas component, which is traced by the H2 0-0 S(1) emission line, shares a similar morphology as that of the [Ne II] line (Fig. 3). The warm molecular gas flux peaks at larger distances from the nucleus and appears more diffuse, closely resembling the CO(3–2) emission. In contrast, the ionized gas traced by [Ne II] is concentrated in more compact regions. Since the spatial resolution at the wavelengths of the two lines is comparable, this difference in morphology reflects an intrinsic difference in the extent of the emitting regions. In addition, we observe a concentration of warm molecular gas at the position of the nucleus, which could be explained by the presence of large amounts of dust (as seen in dust continuum from MIRI by Liu et al., 2023) shielding the molecules from the AGN radiation. The kinematics follows the same rotational pattern observed in [Ne II], with some deviations from a pure rotating disk. In particular, the NW side of the H2 velocity map shows lower blueshifted velocities compared to the [Ne II] velocity. On the opposite side, a deviation from rotation is observed in the redshifted region, which is co-spatial with an enhancement in velocity dispersion. This region, located 2″ south of the nucleus, exhibits a peak in dispersion reaching 80 km s-1 and appears as a distinct region.
The CO(3-2) moment maps shown in Fig. 3 are in agreement with the H2 S(1) kinematics, that is tracing the galaxy rotating disk. A similar CO velocity field was reported by Liu et al. (2023). This finding, which is consistent with the lack of any molecular outflow in the circumnuclear region (see also Combes et al., 2019), is in contrast to the CO (1-0) wind reported by Gao et al. (2021). Their outflow interpretation was based on velocity residuals from a rotating disk model on larger spatial scales, which resemble those observed in H and are more likely due to non-circular inflowing motions along the bar (see their Fig. 6). Notably, the cold gas traced by the CO(3-2) does not show the -enhanced region south of the nucleus as the H2, although both emission lines map the molecular gas. This suggests that the elevated velocity dispersion in H2 may trace non-circular motions, possibly related to local turbulence or streaming motions induced by the AGN, rather than a genuine outflow. In such an environment, the cold molecular gas traced by CO may not survive due to stronger dissociation or heating, leaving only the warm phase observable in this disturbed region.
The moment maps of the other emission lines listed in Table LABEL:tab:_table_list_emission_lines are shown in the online material, divided into molecular gas, recombination lines, and species with low ( 25 eV), medium ( 25 and 54 eV), and high ( 54 eV) IP. The thresholds are based on the ionization potentials of He I (25 eV) and He II (54 eV). Note that while some lines appear to lack emission in the nuclear region of the galaxy in the moment maps, this is due to low S/N where the spectra are dominated by strong continuum emission, rather than a genuine absence of emission. Grouping such emission lines by IP or gas phase leads to interesting typical patterns. All the H2 lines share kinematic features similar to those of H2 S(1) in Fig. 3, with a rotational disk kinematics and a high- blob to the South of the nucleus. The low-IP lines, as well as the recombination lines, show a velocity gradient consistent with a rotating disk, with an amplitude of 110 km s-1, like [Ne II]. The high- region observed in H2 is not present in all the low-IP and recombination lines. From the flux maps, we observed that these species are located mainly in the ‘Mid-East’ region (Liu et al., 2023). The moment maps of the high-IP species are similar to those of [Ne V] (see Fig. 3). In particular, the donut-shaped structure is well defined in the velocity dispersion maps, although in the Ch4 the spatial resolution is lower and the shape is not well recognizable. The medium-IP lines have intermediate properties compared to the previous two groups: the velocity gradient is similar to that of the low-IP lines, but the velocity dispersion maps show a high-velocity dispersion region in correspondence of the donut shape.
3.2 Velocity channel maps
Fig. 4 presents the velocity channel maps of the [O III] and [Ne V] emission lines, approximately spanning the velocity range from km s-1 to +550 km s-1 around the same systemic velocity. Both ionic species share the same kinematic structures in each velocity bin, although [Ne V] appears to trace regions closer to the outflow axis, as also found by Mingozzi et al. (2019) for high-ionization gas in MAGNUM galaxies. This is consistent with the fact that both trace the outflow but are associated with different IPs (97 eV and 35 eV for [Ne V] and [O III], respectively; see Table LABEL:tab:_table_list_emission_lines). Indeed, [Ne V] emission arises from more internal regions of the outflow, where the AGN radiation field is expected to be stronger. Interestingly, Fig. 3 shows that at projected velocities larger than 100 km s-1, the [Ne V] emission resembles the donut-shaped structure mentioned above. The bulk of the emission progressively shifts away from the nucleus with increasing channel velocity, tracing the lower arm of the donut-shaped structure. Moreover, unlike [O III], the redshifted [Ne V] velocity channels clearly reveal the receding NW ionization cone, which is otherwise highly obscured in the optical due to dust in the galaxy disk. Indeed, as discussed in the following, the NW cone lies behind the disk and suffers stronger extinction than the SE cone (Venturi et al., 2017, 2018; Marconcini et al., 2025b). Thanks to the lower dust attenuation in the mid-IR, the [Ne V] line is an optimal tracer of this obscured outflowing component.
3.3 Resolved diagnostic diagrams for gas excitation
Diagnostic diagrams featuring mid-IR emission line ratios are widely used to investigate the nature of the gas ionizing sources (e.g. Hao et al., 2009; Weaver et al., 2010; Inami et al., 2013; Richardson et al., 2022; Martínez-Paredes et al., 2023; Feltre et al., 2023; Garofali et al., 2024; Mingozzi et al., 2025). In particular, to investigate the gas excitation in the circumnuclear region of NGC 1365, we employed mid-IR line ratios recently proposed by Feltre et al. (2023), who analyzed integrated nuclear spectra of 42 local Seyfert galaxies observed with Spitzer, including NGC 1365. To ensure consistent comparisons between line ratios, we corrected the flux maps of each emission line for differences in FoV, pixel scale, and point spread functions (PSF). In practice, for each diagnostic diagram we adopted the FoV, pixel size, and PSF corresponding to the worst case among the lines involved, as detailed below. Additionally, as detailed in Appendix B, we applied multiplicative correction factors to account for flux discontinuities between the datacubes (see Fig. 11).
In the upper left panel of Fig. 5, we show the contours of line ratio model predictions for AGN (Feltre et al., 2023), SF galaxies (Gutkin et al., 2016), and shocks (Alarie & Morisset, 2019) for the [O IV]m/[Ne II]m and [Ne III]m/[Ne II]m line ratios. For this diagnostic diagram, we adopted the smallest FoV (6.6″ 7.7″, i.e. Ch3), the largest pixel size (0.35″ from Ch4), and convolved all images to the worst PSF (i.e. 1.0″, the one at the reddest wavelength of m). We include all available model grids, applying only a metallicity cut of (see Feltre et al. 2023 for further details on the models). We overlay our data on model predictions, where each pixel is color-coded based on its position in the corresponding diagnostic diagram. This diagram shows that the main ionization source in the circumnuclear region of NGC 1365 can either be shocks or AGN, while no evidence for SF excitation is detected. According to Fig. 5 of Feltre et al. (2023), adding fractional contributions from SF and shock models to an AGN model will shift the points towards the bottom-left in the diagram. Therefore, we classify as shock- or AGN-excited only those spaxels that fall entirely within a single predicted region. Data points located in overlapping or unclassified regions cannot be uniquely attributed to a single excitation mechanism. To better visualize the spatial distribution and morphology of these regions, we produce diagnostic maps using the same color-coding of the diagrams, as shown in the lower panels of Fig. 5.
In the central column of Fig. 5, we show the [Ne III]m/[Ne II]m versus [Ne V]m/[Ne II]m diagnostic diagram. In this case, all lines fall within the same channel. Therefore, only the PSF convolution was required, while the FoV and pixel scale were already consistent across the maps. Similarly to the [O IV] diagram, the data points are separated into shock- and/or AGN-excited models. The [O IV] and [Ne V] excitation maps reveal similar spatial structures: an extended region dominated by AGN and combined shock/AGN excitation, aligned with the disk major axis and delimited on either side by shock-dominated zones. Notably, the eastern shock front coincides with the edge of the circumnuclear ring traced by CO and H (see Fig. 3 and Venturi et al., 2018), which may indicate that the outflow, emerging from the disk, interacts with the material of the circumnuclear ring, producing shocks along its path. In contrast, the centrally elongated feature may reflect direct AGN ionization within the disk.
However, Laor (1998) showed that photoionizing shocks are extremely inefficient in powering the narrow line emission in AGN. Furthermore, as noted by Feltre et al. (2023), the predictions from shock models largely overlap with both AGN and star-formation grids in diagnostic diagrams. For these reasons, in Section 3.5 we consider only AGN and SF grids to model the optical and mid-IR emission lines.
Exploiting optical MUSE data, we compare the mid-IR diagnostic diagrams with the optical [S II] BPT diagram (Kewley et al., 2006, 2013) shown in the right column of Fig. 5. Moreover, we also considered the [N II] and [O I] BPT diagrams, finding similar results (see Fig. 5 in Venturi et al., 2018 and Appendix D3 in Mingozzi et al., 2019 for these diagrams at large scale). As found by Venturi et al. (2018), we notice that the circumnuclear SF ring seen in H and in CO (see Fig. 1) is not fully dominated by SF in the BPT diagrams. Instead, as visible from the [S II] spatial distribution, certain regions in the SE of the nucleus show signatures of AGN excitation. Venturi et al. (2018) conclude that this is likely due to the superposition along the line of sight of the AGN-ionized cone and the SF ring. The main difference between the optical BPT and the mid-IR diagnostic diagrams lies in their sensitivity to trace the true ionization source in dusty environments. In the nuclear region, the optical BPT appears to be dominated by star formation due to strong dust obscuration, as indicated by the Balmer decrement (see Appendix D and Fig. 12). In contrast, mid-IR diagnostics can penetrate the dust screen and clearly reveal AGN ionization at the nucleus. This explains why, in the optical, AGN ionization cones are only visible on larger spatial scales, as previously shown in Venturi et al. (2018) and Mingozzi et al. (2019). Another notable difference is that the region classified as shock-dominated in the mid-IR diagrams appears to be excited by the AGN radiation according to the optical diagnostic diagram. This discrepancy likely reflects the different depths probed by optical and mid-IR emission, with mid-IR lines tracing higher ionization regions deeper into the disk, where shocks from the interaction between the outflow and the disk are expected to occur.
3.4 Electron density estimation
One of the main advantages of MIRI is the availability of mid-IR density-sensitive line ratios, such as [Ne V]/[Ne V]. These ratios are sensitive to a higher density regime with respect to the commonly used optical [S II]/[S II] ratio, due to the different ionization energies and critical densities of the lines involved. Recent studies used IR tracers to explore the gas density in local Seyfert galaxies and found densities in the range log(ne/cm-3) = 3–5 (Hermosa Muñoz et al., 2024a, b; Zhang et al., 2024; Ceci et al., 2025; Ramos Almeida et al., 2025).
In the top panel of Fig. 6, we show the electron density map derived from the [Ne V]/[Ne V] ratio using PyNeb (Luridiana et al., 2015), assuming an electron temperature K (following typical values adopted for ionized gas in AGN narrow-line regions; Osterbrock & Ferland, 2006). Across the FoV, we estimate a median electron density of log(/cm-3) = (2.7 0.5), with a peak of log(/cm-3)3.5 at the nucleus. The density cannot be computed in some spaxels where the [Ne V] line ratio exceeds the lower density limit (LDL) of 1.2. In these spaxels, we can assume an upper limit for the density of log(ne/cm-3)2. In the remaining regions, the morphology of the density map appears to follow the conical structure of the outflow, as traced by the [Ne V] emission line (see Fig. 3).
Additionally, as shown in Appendix D, we computed the electron density map from the [S II] doublet observed in the MUSE data (see Venturi et al., 2018 and Mingozzi et al., 2019 for the maps in the entire MUSE FoV). Both the circumnuclear ring and the outflow region are characterized by peaks in electron density, with values reaching log(/cm-3)2.6–2.8. This spatial distribution is consistent with the findings of Venturi et al. (2018) and Kakkad et al. (2018), who reported enhanced electron densities in similar structures. For a quantitative comparison between the optical and mid-IR tracers, we measured the electron density from integrated spectra extracted within a 1.5″ ( 140 pc) radius aperture centered on the outflow region (see Fig. 6), using both the mid-IR and optical tracers. From the mid-IR [Ne V] line ratio, we estimate an electron density of (750440) cm-3, while from the [S II] lines we find a consistent value of (35040) cm-3. These values are illustrated in the bottom panel of Fig. 6, where we plot the theoretical [S II] and [Ne V] line ratios as functions of electron density using PyNeb (Luridiana et al., 2015), assuming an electron temperature K.
Based on our analysis, the gas density derived from the mid-IR diagnostics is 0.3 dex higher than the one derived from the optical, although consistent within the error. This difference is in agreement with the findings of Ramos Almeida et al. (2025), who reported similar, or even larger, offsets in a sample of five Type-2 quasars. Following Ramos Almeida et al. (2025), we also attribute this discrepancy to the different physical conditions probed by the two tracers: optical lines such as [S II] are more sensitive to diffuse, lower-density gas, while mid-IR fine-structure lines trace denser and more obscured regions, typically located closer to the AGN and less affected by dust extinction.
Our attempt to compute from the [Ar V]/[Ar V] ratio within the 1.5″-radius aperture interestingly yielded a ratio of 2.1, which is above the [Ar V] LDL value of 1.8 (corresponding to an upper limit on log(ne/cm-3) density similar to the [Ne V] one, 2). More importantly, even when integrating the [Ne V] flux across the entire MIRI FoV, we do not observe line ratios falling in the low-density limit (LDL), contrary to what was found by Dudik et al. (2007) using Spitzer/IRS data. This further supports the idea that their higher [Ne V]/[Ne V] ratios - and apparent detection of LDL conditions - may originate from observational limitations. In particular, the higher sensitivity of JWST/MIRI MRS compared to Spitzer/IRS may have allowed for a more effective detection of low surface brightness emission that could have remained undetected in previous studies. A direct comparison with the Dudik et al. (2007) measurements and observational setup is presented in Appendix E.
3.5 Photoionization modeling with HOMERUN
We model the full set of optical to mid-IR emission using state-of-the-art photoionization models. Specifically, we employed the Highly Optimized Multi-cloud Emission-line Ratios Using photo-ionizatioN (HOMERUN) modeling framework developed by Marconi et al. (2024).
3.5.1 Model grids of AGN and SF
Given the superposition along the line of sight of the AGN-ionized cone and the circumnuclear SF ring, in the HOMERUN fit we combined two suites of models: one reproducing the emission of dust-free gas ionized by the AGN, and another accounting for the SF contribution, modeled as emission from dusty nebulae around H ii regions and including both dust depletion and dust physics. Both components were assumed to have the same metallicity; however, while we adopted dust-free models for the AGN component, we included dust in the SF models and corrected gas-phase metallicities using the depletion factors listed in Table 7.8 of Hazy 1 (Cloudy v23.1). This choice was motivated by the optical extinction map (Appendix D), which shows clear dust structures associated with the SF ring, but not with the outflowing AGN-ionized gas. Furthermore, the AGN-ionized component exhibits iron coronal lines, indicating that at least part of the dust has been destroyed, releasing iron into the gas phase.
Specifically, the shape of the AGN ionizing radiation field is described by a power law with UV slope , an exponential cutoff exp(-h/k TMax) and the X-ray component slope of =-1.0 linked to the UV through the . We have computed models with following combinations of log(TMax/K) = 4.0, 4.5, 5.0, 5.5, 6.0, 6.5, 7.0 and .
As described in Marconi et al. (2024), the incident spectra for the SF model grids are stellar population models from BPASS v2.3 (Byrne et al., 2022; Stanway & Eldridge, 2018) including binary stellar evolution. These models use a Kroupa (2001) initial mass function and an upper mass cut-off of 300 M⊙. We select models with solar value for the stellar metallicity log(Zstar)=-1.7 and ages in log(age/Myr) of 6.0, 6.4, 6.6, 7.1, 8.8.
The other model parameters are identical for the two suites of models (AGN and SF), namely the ionization parameter log(U) (i.e., the number of ionizing photons compared to that of atoms of neutral hydrogen) ranges from to in step of 0.5 and the hydrogen density of the gas, log(nH / cm-3) from 0 to 7 in steps of 1. For the first iteration HOMERUN spans metallicities log(Z/Z⊙) from -1.0 to 0.4 in steps of 0.2. After this first iteration, HOMERUN performs a final refinement in steps 0.02 on predictions interpolated across the finer grid.
3.5.2 HOMERUN modelling approach
The HOMERUN models the line emission with a weighted combination of multiple CLOUDY single-cloud photoionization models (Ferland et al., 1998), iterating over different gas metallicities and incident radiation fields. The novelty of this method is that the weights assigned to the individual single-cloud models are treated as free parameters during the fitting process.
As explained in detail in Marconi et al. (2024), the fitting process starts by selecting a grid of CLOUDY models, defined by hydrogen density () and ionization parameter () for a fixed ionizing spectrum (), gas phase oxygen abundance (), and elemental abundance ratios (). For each grid, the code searches for the best-fitting multi-cloud model by combining individual single-cloud templates. This is achieved by solving a Non-Negative Least Squares (NNLS) problem, where the weights of each cloud are constrained to be non-negative. The procedure is conceptually analogous to stellar population fitting methods (e.g., pPXF; Cappellari & Emsellem 2004), where a galaxy spectrum is reproduced as a linear combination of stellar templates with positive weights. No regularization is applied in the minimization, allowing for full flexibility in the weight distribution. The process is repeated across all available grids, spanning different ionizing continua, metallicities, and abundance scalings. The optimal solution is identified as the one minimizing the loss function , which quantifies the deviation between model predictions and observed emission-line fluxes. This loss function is effectively a reduced statistic, so values indicate a statistically good fit.
When exploring different metallicities, elemental abundances were scaled from the solar photospheric values of Asplund et al. (2021), except for carbon and nitrogen, which were rescaled following the prescriptions of Nicholls et al. (2017). During the fit, all elemental abundances except oxygen were allowed to vary from their initial values. Scaling factors were used for the AGN and SF components. Since the AGN component was assumed to be dust-free while the SF component includes dust, refractory elements can have different scaling factors (due to depletion), whereas non-refractory elements share the same scaling factor in both components.
3.5.3 HOMERUN fitting results
We fit the fluxes of 60 ionized and neutral atomic emission lines measured from the MIRI and MUSE spectra (32 from MIRI and 28 from MUSE, see Figs. 2 and 9) extracted within a 1.5″ radius aperture centered on the outflow region (see Fig. 6). This aperture was chosen as the largest circular region that is fully covered by all MIRI channels, ensuring the detection of the maximum number of emission lines. The unique approach of HOMERUN enables us to reproduce at the same time 60 different emission lines in different ionisation stages, with the majority of them reproduced with better than 15 accuracy. The data are best matched by a SF component with mean = 3.3, / cm-3) = 1.3, and = -3.3 and an AGN component with mean = 1.4, / cm-3) = 2.9, and = -1.4 (see Fig. 7). Being results of ionized gas region, in the following we will consider .
We find that the SF component is more dust-attenuated than the outflow, in agreement with the extinction map obtained from the Balmer decrement (see Fig. 12) where the higher AV values trace the circumnuclear ring and the nucleus. Interestingly, the difference in AV between SF and AGN components is in line with the median values found in MAGNUM galaxies dividing the emission in disk and outflow (Mingozzi et al., 2019). In Section 3.4, we computed the electron density from the mid-IR [Ne V] and the optical [S II] line doublet ratios within the same 1.5″ radius aperture. The density derived from [Ne V], /cm-3) 2.87, is consistent with the average value of the AGN model components, /cm-3) = 2.9. This is expected since HOMERUN attributes the entire emission of [Ne V] to the AGN (see Table 4), as stars may not produce hard-enough radiation to account for the emission of this transition. Conversely, the [S II] lines yield a relatively higher density of log(ne/cm-3) 2.5 within the 1.5″ radius aperture, compared to log(ne/cm-3) 1.3 from HOMERUN. This is because the average density of the SF component from HOMERUN is lower than that from the [S II] doublet and HOMERUN requires a contribution of more than 70 from the SF component to reproduce the [S II] emission (see Table 3). This highlights the strength of our decomposition method: while classical [S II] diagnostics measure a global electron density, HOMERUN disentangles the different ionizing sources and shows that the SF component alone has a much lower density. Finally, the derived ionization parameters are log(U) = -1.4 for the AGN component and a lower log(U) of -3.3 for the SF component, both in line with values found in H II regions (e.g. Kewley et al., 2001; Dopita et al., 2006) and AGN narrow-line regions (e.g. Groves et al., 2004). This supports the effectiveness of our decomposition in isolating distinct ionization regimes.
Figure 7 shows the results of the fit. The left panel shows the behavior of the loss function as a function of oxygen abundance. The minimum of the curve is found for the minimum loss function of = 0.6, corresponding to 12+log(O/H) = 8.91. The left panels represent the weights of the single-cloud models with different U and NH which were derived for the best fit, for the AGN and SF components in the upper and lower panel, respectively. Using the weights, it is possible to compute the average density and ionization parameter of the clouds, which are identified by the black filled diamond. Only a small number of single-cloud models have non-zero weights, and the AGN and SF components contribute differently across the various lines, depending on the ionization state. In particular, as indicated in Tables 3 and 4, high-ionization lines (e.g., Ne V, O III) are predominantly reproduced by the AGN component, while lower-ionization lines (e.g., N II, Cl II, Ar II) are mainly explained by the SF component. Interestingly, the H luminosity is produced more than 70 by AGN excitation.
Using the same procedure, we fitted with HOMERUN the emission lines in the spectra extracted from the regions marked in Fig. 1 (see Figs. 2 and 9). In the outflow-dominated region (green spectrum), we find for the SF component = 2.8, / cm-3) = 0.9, and = -3.4, and for the AGN component = 1.4, / cm-3) = 3.3, and = -1.4. In the stellar disk region (blue spectrum), we find for the SF component = 2.9, / cm-3) = 1.4, and = -3.3, and for the AGN component = 1.5, / cm-3) = 5.5, and = -1.5. These values are consistent with those obtained from the spectrum extracted within a 1.5″ radius aperture (see Fig. 6). Moreover, SF emission accounts for 64 of the H luminosity in the outflow region and 88 in the stellar disk region.
3.5.4 Ionized gas mass from HOMERUN
A key outcome of the HOMERUN fitting procedure is the ability to relate the luminosity of an emission line to the mass of the ionized gas associated. Since the fit is based on a physically consistent combination of AGN and SF single-cloud models, it allows us to disentangle the contributions from the different ionizing sources and connect the observed emission to intrinsic gas properties. In this way, under the assumption that the outflowing gas is fully ionized by the AGN, we avoided including the contribution from systemic or SF-related components, ensuring that the derived masses trace the outflowing gas itself. In particular, for each component, HOMERUN provides both the intrinsic (i.e., extinction-corrected) luminosities (see Tables 3 and 4) and the corresponding ionized gas mass. From the underlying single-cloud CLOUDY models, one obtains for each model a surface brightness in H (i.e. ) and a gas mass surface density (i.e. ). These model outputs define a direct proportionality between H luminosity and ionized gas mass. By rescaling this relation to the total luminosity of a specific emission line (e.g., [Ne V] or [O III]) and taking into account the AGN fraction as determined from the HOMERUN fit, we derive a general expression:
where is a calibration coefficient derived from the HOMERUN AGN component models. This framework provides a robust and physically grounded method to estimate the mass of ionized outflowing gas from spatially integrated line luminosities.
After applying HOMERUN to NGC 1365 emission line fluxes, we found the following ionized gas masses:
where and are the intrinsic line luminosities emitted by the AGN component, and and are the total observed luminosities (i.e. without the HOMERUN separation between AGN and SF). The scaling factors applied to the observed luminosities account for both the fractional contribution of the AGN component and the effect of reddening. The scaling factors of in Equations 2 and 4 refer to extinction-corrected luminosities and can be directly compared with the standard conversion factors adopted in analytical calculations of the ionized gas masses (Equations 10 and 11, see also the discussion in the next section).
We compute an outflow mass of = (109 2) 103 and = (88 9) 103 within the 1.5″ radius aperture. These differences in mass, obtained with HOMERUN, reflect the discrepancy between the observed and model-predicted line ratios. Indeed, the observed ratio between [O III] and [Ne V] fluxes is 0.81 times the model ratio (see Appendix F), which matches the ratio between the [O III]- and [Ne V]-based outflow masses. We emphasize that the scaling relations in Equations 2-5 for the ionized outflow masses are not universally applicable to other AGN. They are derived specifically from the HOMERUN multi-cloud modeling customized for the physical conditions and emission properties of NGC 1365. Therefore, applying these relations to other galaxies requires performing a dedicated HOMERUN fit tailored to the individual characteristics of each galaxy.
3.6 Outflow energetics
In the previous section, with HOMERUN we obtain an outflow mass of = (109 2) 103 and = (88 9) 103 within a 1.5″ radius aperture centered on the outflow region (see Fig. 6 and Section 3.5.3). In addition, following Carniani et al. (2015), in Appendix G we present a detailed derivation of the outflow mass of [Ne V] and [O III]. Overall, with the classic method we find in the same spatial region = (2.6 1.0) 103 and = (20 7) 104 . The large difference between the HOMERUN and classical mass estimates (factors of 40 for [Ne V] and 2 for [O III]) clearly highlights the fundamental limitations of the standard approach. The classical method relies on several strong assumptions, such as considering that all the species are in the ionization state responsible for the observed line (e.g. O2+ for [O III], Ne4+ for [Ne V]), adopting a single average density (from [S II] in the optical and [Ne V] in the mid-IR), and correcting for attenuation using the Balmer decrement. In contrast, HOMERUN does not require assumptions about ionization levels, temperature, or metallicity, and accounts for differential reddening between the two components. The reliability of the more complex HOMERUN approach is demonstrated by the fact that classical method mass estimates differ by a factor of 75 between different tracers, highlighting the inherent inconsistencies of the traditional approach. Indeed, for example, the HOMERUN fitting shows that the fraction of neon in the form of Ne4+ in the best-fit models is only 10, while for O2+ the fraction is significantly higher, around 60 of the total oxygen. This is certainly one of the main factors contributing to the large discrepancy between the two mass estimates. Moreover, it should be noted that [O III] and [Ne V], having different critical densities and luminosities, trace gas from slightly different physical regions within the outflowing clouds, which further contributes to the observed differences in mass estimates beyond the methodological inconsistencies of the classical approach.
We assume that the flux density within each spaxel is constant over time and that the outflow subtends a solid angle . Therefore, we compute the mass outflow rate as , i.e. the amount of ionized mass crossing a distance with velocity . We also calculate the outflow kinetic energy as and the outflow power as .
| Line | Method | ||||
| [104 M⊙] | [10-1 M⊙ yr-1] | [1053 erg] | [1039 erg s-1] | ||
| [Ne V] | standard | 0.26 0.10 | 0.041 0.020 | 0.04 0.02 | 0.20 0.13 |
| HOMERUN | 10.9 0.2 | 1.7 0.2 | 1.6 0.3 | 8 3 | |
| [O III] | standard | 20 7 | 3.8 0.18 | 4 2 | 26 17 |
| HOMERUN | 8.8 0.9 | 1.7 0.3 | 1.9 0.5 | 12 4 |
We infer the [Ne V] and [O III] intrinsic (i.e. de-projected) radial velocities of the outflow (vout) using the 3D multi-cloud kinematic model (Marconcini et al., 2023). This model has demonstrated exceptional power in reproducing outflow features with unprecedented detail, leveraging IFU data from a variety of instruments (Marconcini et al., 2023; Cresci et al., 2023; Perna et al., 2024; Ulivi et al., 2025; Marconcini et al., 2025a, b). To reproduce the outflow features in NGC 1365, we adopted a conical outflow morphology with maximum extension of 5 arcsec ( 470 pc) and position angle333The position angle is measured clockwise from North. of 235∘. The free parameters of the fit are the inclination of the cone axis with respect to the line of sight () and the intrinsic outflow velocity (vout). With we reproduce the line emission on a spaxel-by-spaxel basis, which as a consequence guarantee that the model reproduces the observed moment maps, with discrepancies between observed and modeled moment maps below 5. We found that the best-fit parameters that reproduce the [Ne V] outflow features in NGC 1365 are = 82 6∘, and vout = 390 35 km s-1, consistent with the results in Marconcini et al. (2025b). Fig. 8 shows the 3D representation of this conical outflow. Moreover, we used to fit the [O III] spatially resolved emission assuming the same outflow geometry as for the [Ne V] emission and found = 84 5∘, and vout = 470 40 km s-1. These results are also valid in the region where we extracted the spectrum used for the HOMERUN fitting.
So, adopting R = 284 pc ( 3″, i.e. the sky-projected extension of the 1.5″ radius aperture), we obtain the results of the outflow energetics reported in Table 1. From HOMERUN, we calculate consistent mass outflow rates for [Ne V] and [O III], namely = M⊙ yr-1 and = M⊙ yr-1, respectively. Instead, with the classic approach the [O III]-derived outflow rate is two orders of magnitude larger than that from [Ne V]. This discrepancy is primarily driven by the underestimation of the [Ne V] outflow mass in the classical approach. Specifically, the classical [Ne V] outflow mass is 75 times smaller than the [O III] estimate and 40 times smaller than the HOMERUN value, while the classical [O III] mass is only a factor of 2 higher than the HOMERUN mass. This asymmetric bias clearly demonstrates that the limitations of the standard method predominantly affects high-ionization tracers like [Ne V], leading to severely underestimated masses and consequently unreliable mass outflow rate calculations.
Our mass outflow rate estimate derived with +HOMERUN models is about one order of magnitude larger than the total ionized outflow rate reported by Venturi et al. (2018), which was derived from the broad H component in a shell-like region of the SE cone, extending from 1″ to 6″ from the nucleus. This large discrepancy clearly highlights how significantly the adopted methodology can affect the inferred outflow properties.
4 Conclusions
In this work, we presented JWST/MIRI MRS observations of the active galaxy NGC 1365 as part of the Mid-IR Activity of Circumnuclear Line Emission (MIRACLE) program. We analyzed the spatially resolved gas properties and compared them with optical and mm observations from MUSE and ALMA, respectively. We traced both the ionized atomic and the warm and cold molecular gas phases in the circumnuclear region ( 0.90.9 kpc2) of the NGC 1365 galaxy, identifying more than 40 mid-IR emission lines. Our main results are the following:
We find that the mid-IR, optical, and mm emission lines consistently trace different phases of the gas with coherent internal kinematics within each phase (see Fig. 3 and Section 3.1). The cold molecular gas traced by the CO(3-2) transition shares the same ordered kinematics of the stellar galaxy disk (Venturi et al., 2018), with no evidence for outflowing material. The same conclusion holds for the warm molecular gas traced by the H2 pure-rotational transitions, with the exception of a peculiar high- unresolved region located southwards of the nucleus.
The ionized gas lines can be separated in two main groups which show opposed morphologies based on their ionization potential (IP). The high-IP ( 54 eV) emission lines trace the bipolar outflow, showing a velocity gradient perpendicular to the disk major axis, consistent with the [O III] kinematics. The velocity dispersion map highlights a donut-shaped structure associated with the approaching side of the outflow which extends up to 2.5″ ( 240 pc). The low-IP ( 25 eV) emission line kinematics resembles the stellar galaxy disk ordered motions, similar to the H2, with clumpy flux peaks in the circumnuclear star-forming ring. Finally, medium-IP lines show intermediate properties, with a kinematics similar to low-IP species but exhibiting the donut-shaped structure typical of the high-IP lines in the velocity dispersion map.
[O III] and [Ne V] velocity channel maps reveal that both outflowing species trace the same kinematic structures, with the [Ne V] appearing more collimated on the outflow axis, as expected from their different IPs (Fig. 4 and Section 3.2). The [Ne V] emission also reveals the NW receding counterpart of the ionization cone, which is undetected in [O III] due to large extinction caused by the galaxy disk at optical wavelengths.
The mid-IR emission-line ratios indicate a composite excitation scenario, with contribution from both AGN photoionization and shocks (Fig. 5 and Section 3.3). The AGN ionization seems to extend along the disk major axis, while in the other direction we observe shock-dominated zones, suggesting an interaction between the outflow and the disk material. Notably, by penetrating the dust in the central regions of the galaxy, these mid-IR diagnostic diagrams reveal for the first time the role of the AGN in ionizing the gas in the nuclear region, differently from the optical diagnostic diagrams that mostly reveal excitation due to SF.
By exploiting the innovative photoionization modeling code HOMERUN, we derived the physical properties of the ionized outflowing gas, providing a self-consistent and physically motivated estimate of the outflow mass (Section 3.5). We obtain = (109 2) 103 and = (88 9) 103 . These estimates differ significantly from those derived with the standard method, which are 40 times lower ([Ne V]) and 2 times higher ([O III]) than those obtained from HOMERUN. This discrepancy reflects the strong assumptions in the classical approach regarding ionization state, average density, and extinction correction.
We find a consistent outflow mass rate of 0.17 M⊙ yr-1 from [Ne V] and [O III] (Section 3.6). Our estimate is about one order of magnitude higher than the value reported in the literature, clearly illustrating how different methods can significantly impact the inferred outflow properties. To derive the intrinsic outflow velocities, we used the 3D kinematic model, assuming a conical geometry. The best-fit solutions provide consistent inclination angles ( 82-84∘) and outflow velocities (vout 390-470 km s-1) for both [Ne V] and [O III], reinforcing the reliability of our kinematic and energetic estimates.
This work highlights the critical need for self-consistent, multi-line photoionization and kinematical modeling to properly capture the complex structure and energetics of AGN-driven outflows - a key step toward understanding their impact on galaxy evolution. We fully exploited the synergy between cutting-edge IFU facilities, JWST/MIRI and VLT/MUSE, to model the ionized gas phase properties across a broad ionization range (from 10 eV to 130 eV), using a total of 60 emission lines. For the first time, we combined detailed 3D kinematic modeling with and photoionization fitting with HOMERUN to derive the intrinsic structure and physical conditions of the outflow in a fully self-consistent framework. This unprecedented approach enabled a coherent and physically motivated characterization of the outflow’s spatial structure, ionization stratification, and energetics. Our results show that classical methods, based on simplified assumptions and a limited set of tracers, significantly underestimate the outflow energetics - and hence the strength of AGN feedback.
Acknowledgements.
We are greatful to S. Charlot for kindly providing the predictions for the SF line-emission models. EB, FB, and GC acknowledge financial support from INAF under the Large Grant 2022 “The metal circle: a new sharp view of the baryon cycle up to Cosmic Dawn with the latest generation IFU facilities” and the GO grant 2024 “A JWST/MIRI MIRACLE: Mid-IR Activity of Circumnuclear Line Emission”. IL, FB, and AM acknowledge support from PRIN-MUR project “PROMETEUS” financed by the European Union - Next Generation EU, Mission 4 Component 1 CUP B53D23004750006. EB acknowledges INAF funding through the “Ricerca Fondamentale 2024” program (mini-grant 1.05.24.07.01). IL and AM acknowledges support from PRIN-MUR project “PROMETEUS” financed by the European Union - Next Generation EU, Mission 4 Component 1 CUP B53D23004750006. AM, MG, IL and CM acknowledge INAF funding through the “Ricerca Fondamentale 2023” program (mini-grant 1.05.23.04.01). FS acknowledges financial support from the PRIN MUR 2022 2022TKPB2P - BIG-z, Ricerca Fondamentale INAF 2023 Data Analysis grant 1.05.23.03.04 “ARCHIE ARchive Cosmic HI & ISM Evolution”, Ricerca Fondamentale INAF 2024 under project 1.05.24.07.01 MINI-GRANTS RSN1 ”ECHOS”, Bando Finanziamento ASI CI-UCO-DSR-2022-43 project “IBISCO: feedback and obscuration in local AGN”. GS acknowledges the project PRIN-MUR 2020 MUR BEYOND-2p (“Astrochemistry beyond the second period elements”, Prot. 2020AFB3FX), the PRIN MUR 2022 FOSSILS (“Chemical origins: linking the fossil composition of the Solar System with the chemistry of protoplanetary disks”, Prot. 2022JC2Y93), the project ASI-Astrobiologia 2023 MIGLIORA (“Modeling Chemical Complexity”, F83C23000800005), the INAF-GO 2023 fundings PROTOSKA (“Exploiting ALMA data to study planet forming disks: preparing the advent of SKA”, C13C23000770005) and the INAF-Minigrant 2023 TRIESTE (“TRacing the chemIcal hEritage of our originS: from proTostars to planEts”; PI: G. Sabatini). SC and GV acknowledge support by European Union’s HE ERC Starting Grant No. 101040227 - WINGS. JF acknowledges financial support from CONAHCyT, project number CF-2023-G100, and UNAM-DGAPA-PAPIIT IN111620 grant, Mexico. AVG acknowledges support from the Spanish grant PID2022-138560NB-I00, funded by MCIN/AEI/10.13039/501100011033/FEDER, EU. MM is thankful for support from the European Space Agency (ESA). EH gratefully acknowledges the hospitality of the IAC, where part of this work was carried out during a long research visit. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. The specific observations analyzed can be accessed via doi: https://doi.org/10.17909/b4w1-hk44. These observations are associated with program 6138. The authors acknowledge the team led by coPIs C. Marconcini and A. Feltre for developing their observing program with a zero-exclusive-access period. This paper makes use of the following ALMA data: ADS/JAO.ALMA2016.1.00296.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.References
- Alarie & Morisset (2019) Alarie, A. & Morisset, C. 2019, Rev. Mexicana Astron. Astrofis., 55, 377
- Alonso-Herrero et al. (2012) Alonso-Herrero, A., Sánchez-Portal, M., Ramos Almeida, C., et al. 2012, MNRAS, 425, 311
- Anand et al. (2021) Anand, G. S., Lee, J. C., Van Dyk, S. D., et al. 2021, MNRAS, 501, 3621
- Argyriou et al. (2023) Argyriou, I., Glasse, A., Law, D. R., et al. 2023, A&A, 675, A111
- Argyriou et al. (2020) Argyriou, I., Wells, M., Glasse, A., et al. 2020, A&A, 641, A150
- Asplund et al. (2021) Asplund, M., Amarsi, A. M., & Grevesse, N. 2021, A&A, 653, A141
- Asplund et al. (2009) Asplund, M., Grevesse, N., Sauval, A. J., & Scott, P. 2009, ARA&A, 47, 481
- Audibert et al. (2023) Audibert, A., Ramos Almeida, C., García-Burillo, S., et al. 2023, A&A, 671, L12
- Bacon et al. (2010) Bacon, R., Accardo, M., Adjali, L., et al. 2010, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 7735, Ground-based and Airborne Instrumentation for Astronomy III, ed. I. S. McLean, S. K. Ramsay, & H. Takami, 773508
- Baldwin et al. (1981) Baldwin, J. A., Phillips, M. M., & Terlevich, R. 1981, PASP, 93, 5
- Belli et al. (2024) Belli, S., Park, M., Davies, R. L., et al. 2024, Nature, 630, 54
- Boroson & Green (1992) Boroson, T. A. & Green, R. F. 1992, ApJS, 80, 109
- Braito et al. (2014) Braito, V., Reeves, J. N., Gofford, J., et al. 2014, ApJ, 795, 87
- Bureau et al. (1996) Bureau, M., Mould, J. R., & Staveley-Smith, L. 1996, ApJ, 463, 60
- Bushouse et al. (2022) Bushouse, H., Eisenhamer, J., Dencheva, N., et al. 2022, spacetelescope/jwst: JWST 1.6.2
- Byrne et al. (2022) Byrne, C. M., Stanway, E. R., Eldridge, J. J., McSwiney, L., & Townsend, O. T. 2022, MNRAS, 512, 5329
- Cano-Díaz et al. (2012) Cano-Díaz, M., Maiolino, R., Marconi, A., et al. 2012, A&A, 537, L8
- Cappellari & Copin (2003) Cappellari, M. & Copin, Y. 2003, MNRAS, 342, 345
- Cappellari & Emsellem (2004) Cappellari, M. & Emsellem, E. 2004, PASP, 116, 138
- Carniani et al. (2015) Carniani, S., Marconi, A., Maiolino, R., et al. 2015, A&A, 580, A102
- CASA Team et al. (2022) CASA Team, Bean, B., Bhatnagar, S., et al. 2022, PASP, 134, 114501
- Ceci et al. (2025) Ceci, M., Cresci, G., Arribas, S., et al. 2025, A&A, 695, A116
- Cicone et al. (2018) Cicone, C., Brusa, M., Ramos Almeida, C., et al. 2018, Nature Astronomy, 2, 176
- Combes (2017) Combes, F. 2017, Frontiers in Astronomy and Space Sciences, 4, 10
- Combes et al. (2019) Combes, F., García-Burillo, S., Audibert, A., et al. 2019, A&A, 623, A79
- Cresci & Maiolino (2018) Cresci, G. & Maiolino, R. 2018, Nature Astronomy, 2, 179
- Cresci et al. (2015) Cresci, G., Marconi, A., Zibetti, S., et al. 2015, A&A, 582, A63
- Cresci et al. (2023) Cresci, G., Tozzi, G., Perna, M., et al. 2023, A&A, 672, A128
- Davies et al. (2024) Davies, R., Shimizu, T., Pereira-Santaella, M., et al. 2024, A&A, 689, A263
- de Vaucouleurs et al. (1991) de Vaucouleurs, G., de Vaucouleurs, A., Corwin, Jr., H. G., et al. 1991, Third Reference Catalogue of Bright Galaxies
- Decleir et al. (2022) Decleir, M., Gordon, K. D., Andrews, J. E., et al. 2022, ApJ, 930, 15
- Dopita et al. (2006) Dopita, M. A., Fischera, J., Sutherland, R. S., et al. 2006, ApJS, 167, 177
- Dudik et al. (2007) Dudik, R. P., Weingartner, J. C., Satyapal, S., et al. 2007, ApJ, 664, 71
- Fabian (2012) Fabian, A. C. 2012, ARA&A, 50, 455
- Feltre et al. (2023) Feltre, A., Gruppioni, C., Marchetti, L., et al. 2023, A&A, 675, A74
- Ferland et al. (1998) Ferland, G. J., Korista, K. T., Verner, D. A., et al. 1998, PASP, 110, 761
- Feruglio et al. (2010) Feruglio, C., Maiolino, R., Piconcelli, E., et al. 2010, A&A, 518, L155
- Fiore et al. (2017) Fiore, F., Feruglio, C., Shankar, F., et al. 2017, A&A, 601, A143
- Fitzpatrick et al. (2019) Fitzpatrick, E. L., Massa, D., Gordon, K. D., Bohlin, R., & Clayton, G. C. 2019, ApJ, 886, 108
- Fluetsch et al. (2021) Fluetsch, A., Maiolino, R., Carniani, S., et al. 2021, MNRAS, 505, 5753
- Fluetsch et al. (2019) Fluetsch, A., Maiolino, R., Carniani, S., et al. 2019, MNRAS, 483, 4586
- Forbes & Norris (1998) Forbes, D. A. & Norris, R. P. 1998, MNRAS, 300, 757
- Förster Schreiber et al. (2019) Förster Schreiber, N. M., Übler, H., Davies, R. L., et al. 2019, ApJ, 875, 21
- Gao et al. (2021) Gao, Y., Egusa, F., Liu, G., et al. 2021, ApJ, 913, 139
- García-Bernete et al. (2024) García-Bernete, I., Alonso-Herrero, A., Rigopoulou, D., et al. 2024, A&A, 681, L7
- García-Burillo et al. (2019) García-Burillo, S., Combes, F., Ramos Almeida, C., et al. 2019, A&A, 632, A61
- Garofali et al. (2024) Garofali, K., Basu-Zych, A. R., Johnson, B. D., et al. 2024, ApJ, 960, 13
- Gordon (2024) Gordon, K. 2024, The Journal of Open Source Software, 9, 7023
- Gordon et al. (2009) Gordon, K. D., Cartledge, S., & Clayton, G. C. 2009, ApJ, 705, 1320
- Gordon et al. (2023) Gordon, K. D., Clayton, G. C., Decleir, M., et al. 2023, ApJ, 950, 86
- Gordon et al. (2021) Gordon, K. D., Misselt, K. A., Bouwman, J., et al. 2021, ApJ, 916, 33
- Groves et al. (2004) Groves, B. A., Dopita, M. A., & Sutherland, R. S. 2004, ApJS, 153, 75
- Gutkin et al. (2016) Gutkin, J., Charlot, S., & Bruzual, G. 2016, MNRAS, 462, 1757
- Hao et al. (2009) Hao, L., Wu, Y., Charmandaris, V., et al. 2009, ApJ, 704, 1159
- Harrison et al. (2016) Harrison, C. M., Alexander, D. M., Mullaney, J. R., et al. 2016, MNRAS, 456, 1195
- Harrison et al. (2014) Harrison, C. M., Alexander, D. M., Mullaney, J. R., & Swinbank, A. M. 2014, MNRAS, 441, 3306
- Harrison & Ramos Almeida (2024) Harrison, C. M. & Ramos Almeida, C. 2024, Galaxies, 12, 17
- Hatziminaoglou et al. (2015) Hatziminaoglou, E., Zwaan, M., Andreani, P., et al. 2015, The Messenger, 162, 24
- Hermosa Muñoz et al. (2024a) Hermosa Muñoz, L., Alonso-Herrero, A., Labiano, A., et al. 2024a, arXiv e-prints, arXiv:2412.14707
- Hermosa Muñoz et al. (2024b) Hermosa Muñoz, L., Alonso-Herrero, A., Pereira-Santaella, M., et al. 2024b, A&A, 690, A350
- Hill & Zakamska (2014) Hill, M. J. & Zakamska, N. L. 2014, MNRAS, 439, 2701
- Inami et al. (2013) Inami, H., Armus, L., Charmandaris, V., et al. 2013, ApJ, 777, 156
- Jacobs et al. (2009) Jacobs, B. A., Rizzi, L., Tully, R. B., et al. 2009, AJ, 138, 332
- Jones & Jones (1980) Jones, J. E. & Jones, B. J. T. 1980, MNRAS, 191, 685
- Kakkad et al. (2018) Kakkad, D., Groves, B., Dopita, M., et al. 2018, A&A, 618, A6
- Kakkad et al. (2020) Kakkad, D., Mainieri, V., Vietri, G., et al. 2020, A&A, 642, A147
- Kewley et al. (2001) Kewley, L. J., Dopita, M. A., Sutherland, R. S., Heisler, C. A., & Trevena, J. 2001, ApJ, 556, 121
- Kewley et al. (2006) Kewley, L. J., Groves, B., Kauffmann, G., & Heckman, T. 2006, MNRAS, 372, 961
- Kewley et al. (2013) Kewley, L. J., Maier, C., Yabe, K., et al. 2013, ApJ, 774, L10
- Kristen et al. (1997) Kristen, H., Jorsater, S., Lindblad, P. O., & Boksenberg, A. 1997, A&A, 328, 483
- Kroupa (2001) Kroupa, P. 2001, MNRAS, 322, 231
- Labiano et al. (2021) Labiano, A., Argyriou, I., Álvarez-Márquez, J., et al. 2021, A&A, 656, A57
- Laor (1998) Laor, A. 1998, ApJ, 496, L71
- Law et al. (2023) Law, D. R., E. Morrison, J., Argyriou, I., et al. 2023, AJ, 166, 45
- Lee et al. (2023) Lee, J. C., Sandstrom, K. M., Leroy, A. K., et al. 2023, ApJ, 944, L17
- Lee et al. (2022) Lee, J. C., Whitmore, B. C., Thilker, D. A., et al. 2022, ApJS, 258, 10
- Lindblad et al. (1996) Lindblad, P. O., Hjelm, M., Hoegbom, J., et al. 1996, A&AS, 120, 403
- Liu et al. (2023) Liu, D., Schinnerer, E., Cao, Y., et al. 2023, ApJ, 944, L19
- Luridiana et al. (2015) Luridiana, V., Morisset, C., & Shaw, R. A. 2015, A&A, 573, A42
- Lutz et al. (2020) Lutz, D., Sturm, E., Janssen, A., et al. 2020, A&A, 633, A134
- Marasco et al. (2020) Marasco, A., Cresci, G., Nardini, E., et al. 2020, A&A, 644, A15
- Marconcini et al. (2025a) Marconcini, C., Feltre, A., Lamperti, I., et al. 2025a, A&A, 701, A113
- Marconcini et al. (2025b) Marconcini, C., Marconi, A., Cresci, G., et al. 2025b, Nature Astronomy, 9, 907
- Marconcini et al. (2023) Marconcini, C., Marconi, A., Cresci, G., et al. 2023, A&A, 677, A58
- Marconi et al. (2024) Marconi, A., Amiri, A., Feltre, A., et al. 2024, A&A, 689, A78
- Martínez-Paredes et al. (2023) Martínez-Paredes, M., Bruzual, G., Morisset, C., et al. 2023, MNRAS, 525, 2916
- Mingozzi et al. (2019) Mingozzi, M., Cresci, G., Venturi, G., et al. 2019, A&A, 622, A146
- Mingozzi et al. (2025) Mingozzi, M., Garcia del Valle-Espinosa, M., James, B. L., et al. 2025, ApJ, 985, 253
- Morrison et al. (2023) Morrison, J. E., Dicken, D., Argyriou, I., et al. 2023, PASP, 135, 075004
- Mulcahey et al. (2022) Mulcahey, C. R., Leslie, S. K., Jackson, T. M., et al. 2022, A&A, 665, A144
- Nicholls et al. (2017) Nicholls, D. C., Sutherland, R. S., Dopita, M. A., Kewley, L. J., & Groves, B. A. 2017, MNRAS, 466, 4403
- Osterbrock & Ferland (2006) Osterbrock, D. E. & Ferland, G. J. 2006, Astrophysics of gaseous nebulae and active galactic nuclei
- Patapis et al. (2024) Patapis, P., Argyriou, I., Law, D. R., et al. 2024, A&A, 682, A53
- Perna et al. (2024) Perna, M., Arribas, S., Lamperti, I., et al. 2024, A&A, 690, A171
- Perrin et al. (2014) Perrin, M. D., Sivaramakrishnan, A., Lajoie, C.-P., et al. 2014, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 9143, Space Telescopes and Instrumentation 2014: Optical, Infrared, and Millimeter Wave, ed. J. M. Oschmann, Jr., M. Clampin, G. G. Fazio, & H. A. MacEwen, 91433X
- Ramos Almeida et al. (2022) Ramos Almeida, C., Bischetti, M., García-Burillo, S., et al. 2022, A&A, 658, A155
- Ramos Almeida et al. (2025) Ramos Almeida, C., Garcia-Bernete, I., Pereira-Santaella, M., et al. 2025, arXiv e-prints, arXiv:2504.01595
- Richardson et al. (2022) Richardson, C. T., Simpson, C., Polimera, M. S., et al. 2022, ApJ, 927, 165
- Richings & Faucher-Giguère (2018a) Richings, A. J. & Faucher-Giguère, C.-A. 2018a, MNRAS, 478, 3100
- Richings & Faucher-Giguère (2018b) Richings, A. J. & Faucher-Giguère, C.-A. 2018b, MNRAS, 474, 3673
- Rieke et al. (2015) Rieke, G. H., Wright, G. S., Böker, T., et al. 2015, PASP, 127, 584
- Riffel et al. (2020) Riffel, R. A., Zakamska, N. L., & Riffel, R. 2020, MNRAS, 491, 1518
- Risaliti et al. (2005) Risaliti, G., Bianchi, S., Matt, G., et al. 2005, ApJ, 630, L129
- Sakamoto et al. (2007) Sakamoto, K., Ho, P. T. P., Mao, R.-Q., Matsushita, S., & Peck, A. B. 2007, ApJ, 654, 782
- Sanchez (2009) Sanchez, R. Z. 2009, PhD thesis, Rutgers University, New Jersey
- Sandqvist et al. (1995) Sandqvist, A., Joersaeter, S., & Lindblad, P. O. 1995, A&A, 295, 585
- Schinnerer et al. (2023) Schinnerer, E., Emsellem, E., Henshaw, J. D., et al. 2023, ApJ, 944, L15
- Sharp & Bland-Hawthorn (2010) Sharp, R. G. & Bland-Hawthorn, J. 2010, ApJ, 711, 818
- Speranza et al. (2024) Speranza, G., Ramos Almeida, C., Acosta-Pulido, J. A., et al. 2024, A&A, 681, A63
- Stanway & Eldridge (2018) Stanway, E. R. & Eldridge, J. J. 2018, MNRAS, 479, 75
- Stevens et al. (1999) Stevens, I. R., Forbes, D. A., & Norris, R. P. 1999, MNRAS, 306, 479
- Teuben et al. (1986) Teuben, P. J., Sanders, R. H., Atherton, P. D., & van Albada, G. D. 1986, MNRAS, 221, 1
- Tozzi et al. (2021) Tozzi, G., Cresci, G., Marasco, A., et al. 2021, A&A, 648, A99
- Ulivi et al. (2025) Ulivi, L., Perna, M., Lamperti, I., et al. 2025, A&A, 693, A36
- Vazdekis et al. (2016) Vazdekis, A., Koleva, M., Ricciardelli, E., Röck, B., & Falcón-Barroso, J. 2016, MNRAS, 463, 3409
- Veilleux et al. (2005) Veilleux, S., Cecil, G., & Bland-Hawthorn, J. 2005, ARA&A, 43, 769
- Veilleux & Osterbrock (1987) Veilleux, S. & Osterbrock, D. E. 1987, ApJS, 63, 295
- Venturi et al. (2021) Venturi, G., Cresci, G., Marconi, A., et al. 2021, A&A, 648, A17
- Venturi et al. (2017) Venturi, G., Marconi, A., Mingozzi, M., et al. 2017, Frontiers in Astronomy and Space Sciences, 4, 46
- Venturi et al. (2018) Venturi, G., Nardini, E., Marconi, A., et al. 2018, A&A, 619, A74
- Venturi et al. (2023) Venturi, G., Treister, E., Finlez, C., et al. 2023, A&A, 678, A127
- Véron-Cetty & Véron (2006) Véron-Cetty, M. P. & Véron, P. 2006, A&A, 455, 773
- Weaver et al. (2010) Weaver, K. A., Meléndez, M., Mushotzky, R. F., et al. 2010, ApJ, 716, 1151
- Woo et al. (2016) Woo, J.-H., Bae, H.-J., Son, D., & Karouzos, M. 2016, ApJ, 817, 108
- Wootten & Thompson (2009) Wootten, A. & Thompson, A. R. 2009, IEEE Proceedings, 97, 1463
- Wright et al. (2023) Wright, G. S., Rieke, G. H., Glasse, A., et al. 2023, PASP, 135, 048003
- Wright et al. (2015) Wright, G. S., Wright, D., Goodson, G. B., et al. 2015, PASP, 127, 595
- Zhang et al. (2024) Zhang, L., Packham, C., Hicks, E. K. S., et al. 2024, ApJ, 974, 195
- Zubovas & Bourne (2017) Zubovas, K. & Bourne, M. A. 2017, MNRAS, 468, 4956
Appendix A Observations and data reduction
In this appendix, we present the observations and data reduction for MIRI MRS, MUSE, and ALMA data.
A.1 JWST/MIRI
We refer to Marconcini et al. (2025a) for a detailed description of the different steps performed during the data reduction and briefly summarize them here. We downloaded the uncalibrated science and background observations through the Barbara A. Mikulski Archive for Space Telescopes (MAST) portal444https://www.vla.nrao.edu/astro/nvas/avla.shtml. The data reduction process was performed using the JWST Science Calibration Pipeline (Bushouse et al. 2022) version with JWST Calibration Reference Data System context file 1298. We applied all three stages of the pipeline processing, which include CALWEBB_DETECTOR1, CALWEBB_SPEC2, and CALWEBB_SPEC3 (see Argyriou et al. 2020, 2023; Morrison et al. 2023; Patapis et al. 2024). For NGC 1365, we subtracted the background emission from the 2D science images using a background frame generated from our dedicated background observations. Moreover, we performed two residual fringe corrections, both in stage 2 and 3. Additionally, to improve the final datacubes, we adopted the Exponential Modified-Shepard Method (EMSM) weighting function, which has proven to be more efficient in reducing the drizzling effect (Law et al. 2023).
Finally, we applied an astrometric correction to all the datacubes with different RA and Dec for each band, using the ALMA astrometry as reference (details in Appendix B).
A.2 MUSE Wide Field Mode
The data were retrieved from the ESO archive555https://archive.eso.org/cms.html and were already processed automatically using the standard MUSE pipeline (v1.6), with an average PSF FWHM of 0.8″. The final datacube covers a FoV of 64″ 63.4″, with a pixel size of 0.2″ pixel-1. The MUSE FoV covers approximately the central 5.3 5.3 kpc2 of NGC 1365. MUSE resolving power in Wide Field Mode (WFM) varies from 1770 at 4800 Å to 3590 at 9300 Å. Finally, we apply a relative astrometric correction, registering the position of the continuum peak, measured in the line-free 7660-8000 Å wavelength range, to the nucleus position from the ALMA data (see Appendix A.3), as done for JWST/MIRI data (see Appendix A.1). The applied corrections were RA and Dec shifts of (-0.78″, +0.22″).
The MUSE integrated spectra of NGC 1365 extracted from the 0.7″ radius regions shown in the lower-left panel of Fig. 1 are reported in Fig. 9. The green (blue) spectrum is extracted from a region where the outflow (stellar disk) emission is predominant. We detect approximately 30 emission lines (see Table 3) tracing different gas phases, including recombination lines, atomic lines, and lines from ionized gas.
A.3 ALMA
The data include observations with two different configurations with angular resolutions in FWHM of ″ and ″, respectively. In this work, we use the observations with resolution ″ corresponding to 30 pc, which better match the MIRI MRS spatial resolution and provide a maximum recoverable scale of 2.9″. We retrieved the calibrated measurement sets using the dedicated service provided by the European ALMA Regional Center (Hatziminaoglou et al. 2015). Imaging was performed using the Common Astronomy Software Applications (CASA) v6.1.1 (CASA Team et al. 2022). For the CO(3–2) spectral window ( GHz), we subtracted a constant continuum level estimated from the emission line-free channels in the uv plane. The data were cleaned using the tclean task in CASA, with briggs weighting scheme with a robust parameter of 0.5, which resulted in a beam with a FWHM of 0.35″ 0.31″ (beam PA = -71∘), corresponding to 37 pc in physical scale. The final datacube has a spectral channel width of 8 km s-1, a pixel size of 0.05″, a FoV of 25″ in diameter and rms of 1.5 mJy beam-1 per channel.
Finally, we fit the central of the CO(3-2) flux map (see Appendix C.3) with a Gaussian profile to determine the coordinates of the nucleus, taking advantage of the high astrometric accuracy provided by ALMA. We find the nuclear position to be RA = 03h33m36.37s and Dec = -36d08m25.37s, in good agreement with the values reported by Liu et al. (2023). These coordinates are adopted throughout the rest of the paper and used to register the astrometry of MIRI and MUSE observations (see Appendices A.1 and A.2, respectively).
Appendix B MIRI datacubes corrections
For the astrometric registration of the 12 datacubes, we applied separate corrections for each band (SHORT, MEDIUM, and LONG), as each observing setup corresponds to a different grating configuration of the MRS. As discussed in Appendix A.1, the observations are centered on the SE outflow cone rather than the AGN. This positioning causes the source to fall near the edge, or even slightly outside, of the field of view in the Ch1 datacubes (see Fig. 1). To ensure accurate alignment across the dataset, we computed the astrometric corrections using Ch3, where the source is well within the FoV and the spatial resolution remains high, providing a reliable reference frame for registration. We collapsed the datacubes along the spectral dimension and fitted the PSFs with a 2D Gaussian model. Then, we aligned the datacubes to match the PSF peaks with the AGN coordinates provided ALMA data (see Appendix A.3). The applied corrections were RA and Dec shifts of (+0.0086″, -0.2655″), (+0.0256″, -0.2713″), and (-0.0300″, -0.299″) for the SHORT, MEDIUM, and LONG channels, respectively.
To compare flux maps of emission lines from different channels it is crucial to account for the different pixel size (see Appendix A.1). To address this, we rebinned each channel to match the pixel size of Ch4, which has the worst resolution (0.35″). We used the reproject exact function to ensure flux conservation during the resampling process. In Fig. 10 we show the collapsed datacubes before and after the spatial rebinning. The relative flux conservation error is below the 1, so negligible respect the original flux.
MIRI data are known to exhibit flux discontinuities between different sub-channels. This effect could be amplified after the rebinning for the different pixel size of data. Therefore, it is essential to apply a scaling factor to the fluxes when constructing a single spectrum, in addition to account for differences in spectral sampling. To address this, we implemented a stitching procedure between each pair of datacubes by rescaling the flux of the bluer datacube to match the median flux value of the overlapping spectral channels of the redder datacube. We chose to rescale to the reddest datacube (Ch4 LONG) as it has the largest FoV, ensuring consistent scaling in every spaxel. This process yielded a scaling factor matrix for each pair of datacubes, which was then applied to correct all fluxes derived from our data. In Fig. 11 an example of the total spectrum before and after the stitching procedure.
As shown in Law et al. (2023), the average FWHM of the MIRI/MRS PSF varies across the four channels spanning the range FWHM = 0.4″, 0.5″, 0.6″, 0.9″, in Ch1, Ch2, Ch3, Ch4, respectively. This variation must be taken into account when computing ratios between emission lines at different wavelengths. To ensure consistent spatial resolution, we firstly downloaded the MIRI MRS PSF model cube using the WebbPSF tool v.1.4.0 (Perrin et al. 2014), which provide a model cube spatially oversampled of a factor of 4 with respect to the observations (i.e. 0.05 ″/pixel for Ch3). Then we deconvolved each flux map with the PSF corresponding to the wavelength of the emission line, and convolved it with the PSF at the reddest wavelength among the lines involved in the ratios, using the PSF model cubes from WebbPSF. The deconvolution procedure is composed by a wavelength-dependent alignment, rotation and flux normalization between the PSFs and our flux maps, following a similar methodology of Marconcini et al. (2025a).
Appendix C Data analysis
In this appendix, we describe the general methodology employed for the emission line analysis of data from MUSE, ALMA, and MIRI MRS. In particular, we analyzed the MUSE and MIRI data to subtract the continuum and fit the emission lines in the optical and the mid-IR band, using a multi-Gaussian fitting procedure.
C.1 JWST/MIRI MRS emission line fitting
This appendix provides further details on the customized python routine we employed to analyze the MIRI MRS datacubes. We followed a three-step procedure: first, we applied spatial smoothing to enhance the S/N; second, we subtracted the local continuum around the brightest emission lines; finally, we performed spaxel-by-spaxel Gaussian fitting of the residual emission. First, we applied a Gaussian spatial smoothing with a kernel size of 1 pixel to every spatial plane of the datacube. This step improves the S/N of faint extended features, at the cost of a moderate degradation in spatial resolution. Specifically, the effective PSF is broadened by approximately 40 across all wavelengths, due to the fact that the pixel size is comparable to or slightly smaller than the intrinsic PSF at each wavelength. Nevertheless, the resulting resolution remains sufficient for our spatially resolved analysis, as it still allows us to distinguish the relevant morphological and kinematic structures in the data.
| Emission line | IP(1) | Ch | Flux | |
| [m] | [eV] | [ erg s-1 cm-2] | ||
| [Mg V] | 5.61 | 109 | 1A | 1.75 0.05 |
| Pf | 7.46 | - | 1C | 0.962 0.005 |
| [Ne VI] | 7.65 | 126 | 2A | 8.44 0.09 |
| [Ar V] | 7.90 | 60 | 2A | 1.4 0.1 |
| H2 0-0 S(4) | 8.03 | - | 2A | 2.06 0.04 |
| [Ar III] | 8.99 | 28 | 2B | 3.4 0.2 |
| H2 0-0 S(3) | 9.66 | - | 2B | 4.6 0.6 |
| [S IV] | 10.51 | 35 | 2C | 24.0 0.4 |
| H2 0-0 S(2) | 12.28 | - | 3A | 7.5 0.1 |
| Hu | 12.37 | - | 3A | 0.760 0.002 |
| [Ne II] | 12.81 | 23 | 3A | 102.6 0.4 |
| [Ar V] | 13.10 | 60 | 3A | 0.702 0.004 |
| [Ne V] | 14.32 | 97 | 3B | 22.2 0.4 |
| [Cl II] | 14.37 | 13 | 3B | 2.02 0.03 |
| [Ne III] | 15.56 | 41 | 3C | 56.6 0.4 |
| H2 0-0 S(1) | 17.04 | - | 3C | 16.3 0.4 |
| [S III] | 18.71 | 23 | 4A | 50.5 0.2 |
| [Ne V] | 24.32 | 97 | 4C | 26.70 0.04 |
| [O IV] | 25.89 | 55 | 4C | 53.0 2.0 |
and the integrated flux within the FoV of the respective channel, expressed in units of erg s-1 cm-2.
[…]
We then performed the emission line fitting spaxel-by-spaxel, fitting each line in Table LABEL:tab:_table_list_emission_lines independently. Prior to this, we modeled and fitted the local continuum selecting wavelength windows on either side of the line, adjusting the range based on the FWHM of each emission line and carefully avoiding other spectral absorption or emission features. Given the short spectral window of 0.1 m we considered, within which the continuum is featureless, we fitted the continuum with a straight line. Each emission line is modeled with both a single and a double Gaussian component, and then we applied a Kolmogorov-Smirnov (KS) test on the residuals to choose the optimal number of Gaussian components in each spaxel, adopting a threshold p-value of 0.7 (see Marasco et al. 2020; Tozzi et al. 2021; Ceci et al. 2025 for details).
In the mid-IR, dust attenuation is much lower than at optical wavelengths, but it does vary with wavelength and affect emission lines differently, potentially impacting specific line ratios. Unfortunately, in our data the S/N of the Pf and Hu (i.e. H I (6-5) at 7.46 m and H I (7-6) at 12.37 m, respectively) is too low to derive a reliable extinction map directly from the mid-IR spectra. To test the potential impact of dust attenuation, we applied an extinction correction to the mid-IR emission lines using the G23 (Gordon et al. 2009, 2021, 2023; Fitzpatrick et al. 2019; Decleir et al. 2022) attenuation curve, from the dustextinction package777https://github.com/karllark/dustextinction (Gordon 2024), and adopting the maximum value (3) derived from our best-fit photoionization model (see Section 3.5.3). However, we find that the resulting variations in the mid-IR line ratios are within the errors, including those involving emission lines at distant wavelengths (e.g., [Ne V]/[Ne V], see Section 3.4).
C.2 MUSE WFM emission line fitting
MUSE data were analyzed as already described in previous works of the MAGNUM program (Venturi et al. 2018; Mingozzi et al. 2019). Briefly, after a preliminary Voronoi tessellation (Cappellari & Copin 2003) to achieve an average S/N per wavelength channel of at least 70 on the continuum, we applied the penalized PiXel-Fitting (pPXF, Cappellari & Emsellem 2004) code to measure the optical emission lines listed in Table 3. Using the penalized PiXel-Fitting (pPXF, Cappellari & Emsellem 2004) code, we fitted simultaneously the continuum and gas emission in the wavelength range 4800-9200 Å, convolving a linear combination of stellar templates (E-MILES; for details see Vazdekis et al. 2016) with a Gaussian velocity distribution. Each emission line was fitted twice, using one and two Gaussian components, and the KS test was applied spaxel-by-spaxel to determine the minimum number of components required to adequately reproduce the line profiles (see Appendix C.1). Given that NGC 1365 hosts a Seyfert 1.8 nucleus, we included an additional spectral component when fitting the bins within a radius of 10 spaxels from the nuclear position to account for the spatially unresolved broad-line region (BLR). The BLR model consists of two independent components: a broad (¿1000 km s-1) Gaussian component for the broad Balmer lines (i.e. H and H) and a template of broad [Fe II] emission features by Boroson & Green (1992). The nuclear model spectrum was obtained by fitting the integrated spectrum of the central spaxels and, in the subsequent spaxel-by-spaxel fitting, it was kept fixed in shape and velocity while being allowed to scale in flux. We then subtracted the stellar and nuclear emission from the original unbinned data cube by rescaling the fitted continuum - assumed constant within each Voronoi bin - to match the observed continuum flux in each spaxel, before performing the subtraction.
We obtained continuum-subtracted datacubes containing the gas emission lines only. Then, we repeated the fitting procedure applied for MIRI JWST data (see previous section). In this case, we used up to three Gaussian components to reproduce the complex line profiles and adopted a p-value of 0.9.
Finally, we corrected the optical emission-line fluxes for dust attenuation using the Balmer decrement (i.e., H/H), assuming a case-B hydrogen recombination value of 2.86 for this ratio (for an electron temperature and density of Te = K and ne = cm-3, respectively; Osterbrock & Ferland 2006) and adopting the G23 attenuation law. The results, presented in Appendix D, are consistent with previous estimates by Venturi et al. (2018) and Mingozzi et al. (2019). As shown in Fig. 12, the dust attenuation peaks at AV = 3.7 in the circumnuclear ring, where young stellar clusters are embedded (Liu et al. 2023; Schinnerer et al. 2023). High attenuation values are also found at the nuclear position, while we do not find any correlation of AV with the outflow morphology, indicative of its low dust content.
C.3 ALMA emission line fitting
We use the CO(3-2) ALMA observations to derive the kinematics of the cold molecular gas. The continuum was subtracted in the -plane during the data reduction (see Appendix A.3). We fit the CO(3-2) line profile in each spaxel using a single Gaussian profile, which is sufficient to reproduce the symmetric line profile. The resulting moment maps of the cold molecular gas are reported in the bottom panels of Fig. 3.
Appendix D Extinction and electron density from MUSE
In this appendix, we present the spatially resolved maps of visual extinction and electron density in the circumnuclear region of NGC 1365, derived from the MUSE data and zoomed-in FoV matching the MIRI one. In the top panel of Fig. 12, we show the visual extinction computed from the Balmer decrement (H/H), adopting the G23 extinction law. High visual extinction values of A 3.7 mag are found at the nuclear position and in the southeastern part of the FoV, corresponding to the circumnuclear ring, as also observed by Venturi et al. (2018) and Mingozzi et al. (2019).
Moreover, the bottom panel of Fig. 12 shows the electron density of the ionized gas, estimated from the total intensity of all the Gaussian components of the [S II] doublet. In particular, we use the [S II]/[S II] diagnostic line ratio (Osterbrock & Ferland 2006), computed over the same FoV of the Ch4. We assume a typical value for the temperature of ionized gas of Te = 104 K and obtain the density map reported in Fig. 12. We observe that the higher values of ne are following the outflow morphology, as traced by [Ne V] contours flux. In the main text, we use as average density the median value and the standard deviation of ne = (230 120) cm-3 measured across the reported FoV.
Appendix E Density comparison with IRS observations
In this appendix, we compare our spatially resolved JWST/MIRI density measurements with those obtained from integrated spectra by Dudik et al. (2007). Dudik et al. (2007) used the Infrared Spectrograph (IRS) on board Spitzer to compute the [Ne V] ratio in a sample of local AGN. They found that Type 2 AGN show higher line ratios with respect to Type 1, with all Type 2 sources lying above the theoretical LDL threshold, implying log(ne/cm-3)2. In particular, for NGC 1365 they computed a [Ne V]/[Ne V] ratio of 2.40.6, which is larger than our estimation of 1.00.2 obtained as the ratio between the two integrated fluxes in the MIRI FoV. They argued that the [Ne V] emitting region likely originates inside the nuclear torus and that differential infrared extinction due to dust in the obscuring torus may be responsible for the ratio above the LDL. Exploiting the spatially resolved [Ne V] emission in our data and the density values in the nuclear region, we can rule out this interpretation. However, given the presence of some LDL regions in Fig. 6, it is possible that their finding is the result of spatial integration over a wide slit, thus including such regions in the measurement. A more plausible reason is an observational effect. Indeed, in Dudik et al. (2007) the [Ne V] fluxes were obtained using the short-wavelength, high-resolution (SH; 4.7″ 11.3″, = 9.9-19.6 m) and long-wavelength, high-resolution (LH; 11.1″ 22.3″, = 18.7-37.2 m) modules of IRS. As it is clear from the moment maps of the high-IP species in the online material, the extension of [Ne V] emission is detected up to 8″ from the nucleus. This means that the SH module surely lost part of the [Ne V] flux, leading to an artificially enhancement of the [Ne V]/[Ne V] line ratio.
Appendix F HOMERUN emission line fluxes
| Emission line | [Å] | Obs. Luminosity | Model | Mod/Obs | AGN frac. | SF frac. | AGN dered. | SF dered. |
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) |
| H | 4861 | |||||||
| [O III] | 5007 | |||||||
| [N I]∗ | 5199 | |||||||
| He II | 5411 | |||||||
| [Fe VII] | 5721 | |||||||
| [N II] | 5755 | |||||||
| [He I]∗ | 5876 | |||||||
| [Fe VII] | 6087 | |||||||
| [O I] | 6300 | |||||||
| [S III] | 6312 | |||||||
| [O I] | 6364 | |||||||
| [Fe X] | 6375 | |||||||
| H | 6563 | |||||||
| [N II] | 6583 | |||||||
| He I | 6678 | |||||||
| [S II] | 6716 | |||||||
| [S II] | 6731 | |||||||
| [Ar V] | 7006 | |||||||
| [He I]∗ | 7065 | |||||||
| [Ar III] | 7136 | |||||||
| Ca II | 7291 | |||||||
| [O II]∗ | 7323 | |||||||
| [O II]∗ | 7332 | |||||||
| [Ni II] | 7378 | |||||||
| [Ar III] | 7751 | |||||||
| [Cl II] | 8579 | |||||||
| [Fe II] | 8617 | |||||||
| [S III] | 9069 |
| Emission line | [m] | Obs. Luminosity | Model | Mod/Obs | AGN frac. | SF frac. | AGN dered. | SF dered. |
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) |
| [Fe II] | 5.34 | |||||||
| [Fe VIII] | 5.44 | |||||||
| [Mg VII] | 5.50 | |||||||
| [Mg V] | 5.61 | |||||||
| [Ni II] | 6.64 | |||||||
| [Fe II] | 6.72 | |||||||
| [Ar II] | 6.98 | |||||||
| [Na III] | 7.32 | |||||||
| Pa | 7.46 | |||||||
| [Ne VI] | 7.65 | |||||||
| [Fe VII] | 7.81 | |||||||
| [Ar V] | 7.90 | |||||||
| [Na VI] | 8.61 | |||||||
| [Ar III] | 8.99 | |||||||
| [Na IV] | 9.04 | |||||||
| [Fe VII] | 9.53 | |||||||
| [S IV] | 10.51 | |||||||
| Hu | 12.37 | |||||||
| [Ne II] | 12.81 | |||||||
| [Ar V] | 13.10 | |||||||
| [Mg V] | 13.52 | |||||||
| [Ne V] | 14.32 | |||||||
| [Cl II] | 14.37 | |||||||
| [Na VI] | 14.40 | |||||||
| [Ne III] | 15.56 | |||||||
| [Fe II] | 17.94 | |||||||
| [S III] | 18.71 | |||||||
| [Fe VI] | 19.56 | |||||||
| [Fe III] | 22.93 | |||||||
| [Ne V] | 24.32 | |||||||
| [O IV] | 25.89 | |||||||
| [Fe II] | 25.99 |
Appendix G Ionized gas mass estimate
In this appendix we present an analytical mass estimation for the [Ne V] and the [O III] emission lines, in order to compute the gas masses in the 1.5″ radius aperture (shown as a dashed circle in Fig. 12) and compare them with the HOMERUN estimations in Section 3.5.4.
Following Carniani et al. (2015), we start from the luminosity of the [Ne V] line:
with f the filling factor, V the volume occupied by the outflowing ionized gas, the electron density, the density of Ne4+ ions, and the line emissivity. can be written as
We assume , 1.2 (i.e. a 10% number density of He atoms with respect to H atoms), and a solar abundance of [Ne/H] 7.93 (Asplund et al. 2009), to obtain . Assuming a typical temperature (T), from the [Ne V]/[Ne V] line ratio we find an electron density of ne = (750440) cm-3 in the 1.5” radius aperture and estimate a line emissivity of erg s-1 cm3 making use of PyNeb (Luridiana et al. 2015). Finally, Eq. 6 can be rewritten as
assuming a constant electron density in the outflow volume. The gas mass can be expressed as
where the average molecular weight is = 1.27 , if there is a 10% fraction of He atoms, and we have taken into account that
therefore, . Combining this latter with Equation 7, we finally get
Similarly, we compute the [O III] outflow mass. Using a solar abundance of [O/H] 8.69 (Asplund et al. 2009) and ne = (500 270) cm-3 (from the [S II] ratio in the same aperture) - that leads to erg s-1 cm3 - we have
We corrected both the [Ne V] and the [O III] luminosity for dust attenuation using the Balmer decrement from the Hα/Hβ ratio, employing the G23 attenuation law and an intrinsic ratio Hα/Hβ = 2.86 (for Te = 104 K, Osterbrock & Ferland 2006). Assuming the two different densities for the optical and mid-IR species previously computed, we obtain an outflow mass of = (2.6 1.0) 103 and = (20 7) 104 .